# FedFDP: Fairness-Aware Federated Learning with Differential Privacy

### v3

@@ Appendix A Appendix
Table 4: Summary of main notations
A.1 Extra Experiment
Datasets MNIST FashionMNIST CIFAR10
Accuracy Fairness Accuracy Fairness Accuracy Fairness
Table 5: FedFair and baselines: results of test accuracy (%) and fairness ($\Psi$).
Herterogeneity Scalability
Datasets MNIST FashionMNIST CIFAR10
FedAvg 98.67 |||| 7.3e-3 98.37 |||| 1.2e-3 89.41 |||| 3.6e-2 62.03 |||| 1.1e0 60.89 |||| 1.5e0 59.78 |||| 1.5e0
SCAFFLOD 98.45 |||| 5.0e-3 98.51 |||| 2.1e-3 86.39 |||| 3.4e-2 62.21 |||| 9.9e0 59.48 |||| 8.7e0 58.64 |||| 2.3e0
FedProx 98.70 |||| 7.5-3 98.35 |||| 1.2e-3 88.33 |||| 3.5e-2 62.13 |||| 1.2e0 60.33 |||| 2.2e0 61.31 |||| 1.6e0
FedDyn 98.40 |||| 8.0e-3 98.36 |||| 1.8e-3 88.26 |||| 3.5e-2 62.33 |||| 9.4e-1 61.35 |||| 2.6e0 62.52 |||| 1.5e0
q-FFL 98.72 |||| 4.7e-3 98.61 |||| 1.1e-3 89.03 |||| 3.6e-2 63.33 |||| 9.4e-1 60.63 |||| 1.3e0 59.65 |||| 1.5e0
FedFair 98.75 |||| 3.9e-3 98.72 |||| 1.2e-3 89.60 |||| 1.7e-2 62.38 |||| 8.5e-1 61.58 |||| 1.1e0 62.90 |||| 1.5e0
Table 6: FedFair and baselines: results of test accuracy (%) and fairness ($\Psi$) in different heterogeneity and scalability.
Default Settings
We conducted experiments on three computer vision (CV) datasets: MNIST (LeCun et al. 1998), FashionMNIST (Xiao, Rasul, and Vollgraf 2017), and CIFAR10 (Krizhevsky 2009). And we comparised our algorithms to six baselines:
  1. 1.
    FedAvg(McMahan et al. 2017): FedAvg is the most classic algorithm in federated learning. In fact, we can set $\lambda$ to 0, transforming FedFair and FedFDP into the FedAvg algorithm.
  2. 2.
    SCAFFLOD(Karimireddy et al. 2020): SCAFFOLD uses control variates (variance reduction) to correct for the “client-drift” in its local updates.
  3. 3.
    FedProx(Li et al. 2020a): FedProx introduces a proximal term to the objective function to address the issues of systems heterogeneity and statistical heterogeneity in federated learning.
  4. 4.
    FedDyn(Acar et al. 2021): FedDyn adds linear and quadratic penalty terms to the objective function to solve the problem of inconsistency between the minima of the local-device level empirical loss and those of the global empirical loss.
  5. 5.
    ALI-DPFL(Ling et al. 2024): ALI-DPFL dynamically obtains the optimal number of local DPSGD iterations for clients between any two sequential global updates through minimizing the convergence bound.
  6. 6.
    q-FFL(Li et al. 2020b): q-FFL amplifies loss values for clients with poorer performance by assigning them larger update weights, aiming to improve their model performance and achieve a balanced performance across multiple devices.
For all datasets, we adopted a widely used heterogeneous settings (Lin et al. 2020; Li, He, and Song 2021) to 10 clients which controlled by a Dirichlet distribution denoted as $\mathrm{Dir}(\beta)$, and the default value of $\beta=0.1$ (Lin et al. 2020; Wang et al. 2020). We use a 4-layer CNN architecture (McMahan et al. 2017) which consists of two convolutional layers and two fully connected layers as model architecture.
In hyper-parameters, for FedFair and its baseline algorithms, we set learning rate $\eta=0.1$. For FedFDP and its baseline algorithms, we set the $\eta=1.0$, the batch sample ratio $q=0.05$, clipping bound for gradient $C=0.1$, the noise multipliers for gradient $\sigma=2.0$, the privacy budget $\epsilon=3.52$ and the $\delta=1e-5$. In particular, for the loss values in the FedFDP algorithm that additionally require differential privacy processing, we set the clipping bound for loss $C_{l}=2.5$ and the noise multipliers for loss $\sigma_{l}=5.0$.
Extended description of FedFair
@@ 21      $\mathbf{w}_{t+1}^{i}\leftarrow\mathbf{w}_{t}^{i}$ , compute $F_{i}(\mathbf{w}_{t+1}^{i})$;
  • Server (lines 2-2): The server executes $T$ rounds of communication, each round sending the aggregated model $\mathbf{w}_{t}$ and the loss $F(\mathbf{w}_{t})$ to clients.
  • Local (lines 2-2): For each batch $b$, we compute the adaptive learning rate $\eta_{i}$ in lines 2-2. For line 2, we execute gradient descent by using $\eta_{i}$.
Result of FedFair
Furthermore, as the training process advances, the model progressively prioritizes fairness, leading to a reduction in $\eta_{i}$. In comparison to training methods with a fixed learning rate, learning rate decay can accelerate the convergence of gradient-based methods (Zinkevich 2003). Performance and Fairness. Table 5 presents the experimental results of the FedFair algorithm compared to baseline methods. FedFair significantly enhances fairness while maintaining substantial model performance across the MNIST, FashionMNIST, and CIFAR10 datasets, with respective increases of 17.0%, 25.7%, and 9.6%.
Heterogeneity. Table 6 shows that FedFair outperforms baseline methods in accuracy while maintaining significant fairness when faced with diverse data distributions exhibiting varying levels of heterogeneity. Except for the $\mathrm{Dir}(0.5)$ scenario, where it falls slightly short of q-FFL, FedFair demonstrated fairness improvements of 22.0% and 50% over the best baseline in $\mathrm{Dir}(0.1)$ and $\mathrm{Dir}(1.0)$, respectively.
Scalability. Table 6 indicates that FedFair generally surpasses baselines in accuracy and shows fairness improvements of 9.6%, 15.4%, and 0.01% across three client counts.
A.2 Proof of Theorem 1
@@
In our analysis, we define two virtual sequences $\mathbf{v}_{t}=\sum_{i=1}^{N}p_{i}\mathbf{v}_{t}^{i}$ and $\mathbf{w}_{t}=\sum_{i=1}^{N}p_{i}\mathbf{w}_{t}^{i}$, which is motivated by (Stich 2018). Therefore,
$\mathbf{v}_{t+1}$ $=\mathbf{w}_{t}$
$\quad-\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})]$ (18)
Key Lemma
Lemma 1
(Results of one iteration.)Assume Assumption 1-3 hold, we have:
@@
$\Delta_{t+1}\leq\left(1-\mu\eta C_{t}\right)\Delta_{t}+\eta^{2}A$
STEP 1. When $t=1$, the equation $\Delta_{1}\leq v$ holds obviously.
STEP 2. We assume $\Delta_{t}\leq\frac{v}{t}$ holds.
STEP 3.
$\Delta_{t+1}$ $\leq\left(1-\mu\frac{\beta}{t}C_{t}\right)\frac{v}{t}+\frac{\beta^{2}A}{t^{2}}$
$=\frac{t-1}{t^{2}}v+\left(\frac{\beta^{2}A}{t^{2}}-\frac{\mu\beta C_{t}-1}{t^{2}}v\right)$
$\leq\frac{t-1}{t^{2}}v$
$\leq\frac{v}{t+1}$
Then by the $L$-smoothness of $F(\cdot)$, let $\beta=\frac{2}{\mu}$, we get
Proof of Lemma 1
By the Equation (A.2), we get
$\left\|\mathbf{v}_{t+1}-\mathbf{w}^{*}\right\|^{2}$
$=\|\mathbf{w}_{t}-\mathbf{w}^{*}$
$-\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})]\|^{2}$
$=\left\|\mathbf{w}_{t}-\mathbf{w}^{*}\right\|^{2}$
Firstly, we process $\mathcal{A}_{2}$:
where $\frac{1}{\hat{B}^{2}}\triangleq\max_{i}\frac{1}{|\mathcal{B}_{i}|}$.
By the convexity of $\|\cdot\|^{2}$,
taking exception and according to assumption 3:
Now, we obtain the bound for the expectation of $\mathcal{A}_{2}$:
The process of $\mathcal{A}_{1}$ show as below:
$\mathcal{A}_{1}$ $=-2\langle\mathbf{w}_{t}-\mathbf{w}^{*},\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\rangle$
$\quad\underbrace{-2\langle\mathbf{w}_{t}-\mathbf{w}^{*},\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}\sigma C\mathcal{N}(0,\mathbf{I})\rangle}_{\mathcal{C}_{0}}$
$=\mathcal{C}_{0}$
$\quad\underbrace{-2\sum_{i=1}^{N}p_{i}\langle\mathbf{w}_{t}-\mathbf{w}_{t}^{i},\frac{\eta}{|\mathcal{B}_{i}|}\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\rangle}_{\mathcal{C}_{1}}$
$\quad\underbrace{-2\sum_{i=1}^{N}p_{i}\langle\mathbf{w}_{t}^{i}-\mathbf{w}^{*},\frac{\eta}{|\mathcal{B}_{i}|}\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\rangle}_{\mathcal{C}_{2}}$
It’s obvious that $\mathbb{E}[\mathcal{C}_{0}]=0$.
By Cauchy-Schwarz inequality and AM-GM inequality, we have
$=\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}-\mathbf{w}_{t}^{i}\|^{2}+\mathcal{B}_{1}$
So we get
According to Assumption 2, we know that
So we get
$\mathcal{C}_{2}$ $\leq 2\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}\sum_{j=1}^{|\mathcal{B}_{i}|}C_{t}^{i,j}\cdot$
$\quad\left[-\left(F_{i}(\mathbf{w}_{t}^{i},\xi_{j})-F_{i}(\mathbf{w}^{*})\right)-\frac{\mu}{2}\|\mathbf{w}_{t}^{i}-\mathbf{w}^{*}\|^{2}\right]$
$=-2\eta C_{t}\sum_{i=1}^{N}p_{i}\left(F_{i}(\mathbf{w}_{t}^{i})-F^{*}+F^{*}-F_{i}(\mathbf{w}^{*})\right)$
$\quad-\mu\eta C_{t}\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}^{i}-\mathbf{w}^{*}\|^{2}$
Next, we proceed to handle $\mathcal{D}_{1}$.
(from the Assumption 2)
$\geq-\frac{1}{2}\sum_{i=1}^{N}p_{i}\left[\eta\|\nabla F_{i}(\mathbf{w}_{t})\|^{2}+\frac{1}{\eta}\|\mathbf{w}_{t}^{i}-\mathbf{w}_{t}\|^{2}\right]$
$\quad+(F(\mathbf{w}_{t})-F^{*})$
(from the AM-GM inequality)
$\geq-\sum_{i=1}^{N}p_{i}\left[\eta L(F_{i}(\mathbf{w}_{t})-F_{i}^{*})+\frac{1}{2\eta}\|\mathbf{w}_{t}^{i}-\mathbf{w}_{t}\|^{2}\right]$
$\quad+(F(\mathbf{w}_{t})-F^{*})$
(from the L-smooth inference)
where L-smooth inference as show:
Thus, we get
To sum up,
$\quad+\eta^{2}\left(G^{2}C_{t}^{2}+2L\Gamma C_{t}\right),$
and
So we get,
$\mathbb{E}\left[\mathcal{A}_{1}\right]$ $=(1-\mu\eta C_{t})\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}+(1+C_{t})\eta^{2}C_{t}^{2}G^{2}$
$\quad+\eta^{2}\left(2G^{2}C_{t}^{2}+2L\Gamma C_{t}+2\frac{\sigma^{2}C^{2}d}{\hat{B}^{2}}\right)$
All in all, we get
where
$A=G^{2}C_{t}^{3}+3G^{2}C_{t}^{2}+2L\Gamma C_{t}+\frac{2\sigma^{2}C^{2}d}{\hat{B}^{2}}.$
A.3 Proof of Theorem 2
When $t\rightarrow T$ and $T$ is sufficiently large, $\|\nabla F_{i}(\mathbf{w}_{t}^{i},x_{j})\|\rightarrow 0$, $C_{t}^{i,j}=1+\lambda\Delta_{i}^{j}$. Assuming $Q_{0}\leq\Delta_{i}^{j}\leq Q_{1}$, then $1+Q_{0}\lambda\leq C_{t}^{i,j}\leq 1+Q_{1}\lambda$. We rearrange the convergence upper bound in Equation (14) to obtain a function in terms of $\lambda$,

### v4

@@ III Preliminaries
III-A Federated Learning and Fairness
We have defined balanced performance fairness in FL as follows.
Definition 1
(Balanced performance fairness [28]) For trained models $\mathbf{w}_{1}$ and $\mathbf{w}_{2}$, we informally recognize that model $\mathbf{w}_{1}$ provides a more fair solution to the federated learning objective in Equation (1) than model $\mathbf{w}_{2}$ if the performance of model $\mathbf{w}_{1}$ on the $N$ devices is more uniform than the performance of model $\mathbf{w}_{2}$ on the $N$ devices.
III-B Differential Privacy
Definition 2
(Differential Privacy [11]). $(\epsilon,\delta)$-differential privacy is achieved by a randomized mechanism $\mathcal{M}:\mathcal{X}^{n}\rightarrow\mathbb{R}^{d}$, for any two neighboring databases $D_{i},D_{i}^{\prime}\in\mathcal{X}^{n}$ that differ in only a single entry, and $\forall$ $\mathcal{S}\subseteq\text{Range}(\mathcal{R})$:
DPSGD. In gradient-based methods, DPSGD[1] is widely used in privacy preservation. It performs per-sample gradient clipping and noise addition:
  • Firstly, computing the gradient for data sample $\xi_{j}$: $g^{i,j}_{t}=\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})$.
  • Secondly, clipping the gradient of per-sample to get the sensitivity:
    $\hat{g}^{i,j}_{t}=g^{i,j}_{t}/\max\left(1,\frac{\|g^{i,j}_{t}\|}{C}\right),$ (4)
    where $C$ represents the clipping bound.
  • Thirdly, adding Gaussian noise with mean 0 and standard deviation $C\sigma$ to the sum of clipped gradients, where $\sigma$ is the noise multiplier:
    $\tilde{g}^{i}_{t}=\frac{1}{|\mathcal{B}_{i}|}\left(\sum_{j=1}^{|\mathcal{B}_{i}|}\hat{g}^{i,j}_{t}+C\sigma\cdot\mathcal{N}(0,\mathbf{I})\right).$ (5)
IV FedFair: Equipping FL with Fairness
@@ 21      $\mathbf{w}_{t+1}^{i}\leftarrow\mathbf{w}_{t}^{i}$ , compute $F_{i}(\mathbf{w}_{t+1}^{i})$;
$\min_{\mathbf{w}}H(\mathbf{w})=F(\mathbf{w})+\frac{\lambda}{2}\sum_{i=1}^{N}p_{i}\left(F_{i}(\mathbf{w})-F(\mathbf{w})\right)^{2}.$ (6)
where $F(\mathbf{w}^{i}_{t})\triangleq\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}^{i}_{t})$ and $F_{i}(\mathbf{w}^{i}_{t})$ is determined using the training dataset [^1] We use $F(\mathbf{w}_{i}^{t})\leftarrow F(\mathbf{w}_{t})\triangleq\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}_{t}^{i})$, the reason for $F(\mathbf{w}_{t})\neq F(\mathbf{w}_{t},\xi_{test})$ is that the test data $\xi_{test}$ from the server cannot be used for training. from round $t-1$, as delineated in line 1 of Algorithm 1, and the $\lambda$ ($\lambda\geq 0$) is a hyperparameter used to adjust the degree of fairness  [^2] When $\lambda=0$, the algorithm becomes traditional federated learning. When $\lambda\rightarrow\infty$, the algorithm only focus on fairness. However, if $\lambda$ is too large, it will cause very steep gradients which might prevent the model from converging. Fortunately, we theoretically found an optimal value for $\lambda$ in Section VI-B. In addition, as shown in Figure 3, we experimentally confirmed the existence of the optimal $\lambda$.. Consequently, within the context of Equation (7), $F(\mathbf{w}^{i}_{t})$ is considered a constant when the gradient is being computed. So, the gradient of Equation (7) as follows:
$\nabla H_{i}(\mathbf{w}_{t}^{i})$ $=\nabla F_{i}(\mathbf{w}_{t}^{i})+\lambda\Delta_{i}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i})$
$=(1+\lambda\Delta_{i})\nabla F_{i}(\mathbf{w}_{t}^{i}),$ (8)
Figure 2: FedFDP framework.
where $\Delta_{i}=F_{i}(\mathbf{w}_{t}^{i})-F(\mathbf{w}_{t}^{i})$.
The stochastic gradient descent by using Equation (IV) as show:
where $\eta_{i}=\eta(1+\lambda\Delta_{i})$.
Therefore, our approach can be recognized as establishing an dynamic learning rate that allows the model update to achieve both performance and fairness objectives simultaneously. Based on these principles, we introduce the FedFair algorithm, as describe in Algorithm 1:
  • Server (lines 1-1): The server executes $T$ rounds of communication, each round broadcasting the aggregated model $\mathbf{w}_{t}$ and the loss $F(\mathbf{w}_{t})$ to clients.
  • Local (lines 1-1): For each batch $b$, we compute the adaptive learning rate $\eta_{i}$ in lines 1-1. For line 1, we execute gradient descent by using $\eta_{i}$.
V FedFDP: Further Equipping FedFair with Differential Privacy
Next, we discuss the details of these two strategies.
V-A Fair-clipping Strategy for Gradient
In the general DPSGD algorithm [1, 14, 4, 49], per-sample clipping is an essential process for obtaining sensitivity. In order to be compatible with per-sample clipping, initially, we expand our objective function of client $i$ in FedFair, Equation (7), into Equation (10):
Next, we compute the gradient of Equation (10):
where $\Delta_{i}^{j}=F_{i}(\mathbf{w}_{t}^{i},\xi_{j})-F(\mathbf{w}_{t}^{i})$.
To ensure the algorithm is safeguarded by differential privacy, we replace the gradient of the loss function with Equation (11) and reformulate the gradient descent computation for client $i$ as follows:
$\mathbf{w}_{t+1}^{i}$ $=\mathbf{w}_{t}^{i}-\eta\tilde{g}^{i}_{t}$
$=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}\frac{\left(1+\lambda\cdot\Delta_{i}^{j}\right)\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})}{\max\left(1,\frac{\left(1+\lambda\cdot\Delta_{i}^{j}\right)\|\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\|}{C}\right)}$
$\quad+\sigma C\cdot\mathcal{N}(0,\mathbf{I})]$
$=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})],$ (12)
@@ V-A Fair-clipping Strategy for Gradient
V-B Adaptive Clipping Method for Loss
Since calculating $\Delta_{i}^{j}$ requires clients to upload $F_{i}(\mathbf{w}_{t}^{i})$, differential privacy must be incorporated to ensure the algorithm preserves privacy. As shown in lines 2-2 of Algorithm 2, $F_{i}(\mathbf{w}_{t}^{i})$ must be clipped to the interval $[0,C_{l}^{i,t}]$, followed by the addition of Gaussian noise with a mean of 0 and a standard deviation of $\sigma_{l}\cdot C_{l}^{i,t}$ to ensure differential privacy. However, determining an appropriate clipping bound $C_{l}^{i,t}$ for $F_{i}(\mathbf{w}_{t}^{i})$ is challenging. A large $C_{l}^{i,t}$ prevents distortion of the original values but results in larger injected noise.
where
Input: loss function $F(\mathbf{w})$, learning rate $\eta$, noise multiplier for gradient $\sigma$, noise multiplier for loss $\sigma_{l}$, fairness hyperparameter $\lambda$, batch sample ratio $q$, original clipping bound $C$.
Output: the final trained model $\mathbf{w}_{T}$
1 SERVER
2       for $t\in[0,1,\cdots,T-1]$  do
3             for $i\in[1,2,\cdots,N]$ parallel  do
5
6            $\mathbf{w}_{t+1}=\sum_{i=1}^{N}p_{i}\mathbf{w}_{t+1}^{i}$;
7             $F(\mathbf{w}_{t+1})=\sum_{i=1}^{N}p_{i}\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$;
8
9      return $\mathbf{w}_{T}$;
10
11LOCAL
13       Sample randomly a batch $\mathcal{B}_{i}$ with probability $q$;
14       for $j\in[1,2,\cdots,|\mathcal{B}_{i}|]$  do
15             $\Delta_{i}^{j}=F_{i}(\mathbf{w}_{t}^{i},\xi_{j})-F(\mathbf{w}_{t})$;
16             Compute $C^{i,j}_{t}$ by Equation (13);
17             $\tilde{g}_{t}^{i,j}=C_{t}^{i,j}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})$;
18
20
21      Compute clipping bound of loss $C_{l}^{i,t}$ by Equation (14) and (15) ;
22       for $j\in[1,2,\cdots,|\mathcal{B}_{i}|]$  do
23             Compute $F_{i}(\mathbf{w}_{t+1}^{i},\xi_{j})$;
24
25            $f_{i,j}=\min({C_{l}^{i,t}},\max(0,F_{i}(\mathbf{w}_{t+1}^{i},\xi_{j})))$;
26
28       return $\mathbf{w}_{t+1}^{i},\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$;
29
Algorithm 2 FedFDP
To ensure differential privacy during the clipping process, we use the noise-added loss $C_{l}^{i,t}$ as the clipping bound. For each communication round, the clipping bound follows the post-processing property of differential privacy.
V-C Overall Algorithm
  • DPSGD (lines 2-2): In lines 2-2, we compute the fair-clipping bound $C_{t}^{i,j}$ and the per-sample gradient $\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})$ to get the processed gradient $\tilde{g}_{t}^{i,j}$. Then, we add mean 0 noise to the sum of per-sample processed gradient $\tilde{g}_{t}^{i,j}$ in line 2 and execute the gradient decent.
  • DP for local loss (lines 2-2): As described in Section V-B, add noise to the clipped local loss by using adaptive clipping bound $C_{l}^{i,t}$.
VI Finding Optimal Fairness Parameter
VI-A Convergence Analysis
Assumption 1
Assumption 2
Assumption 3
Let $\xi_{j}$ be sampled from the $i$-th device’s local data uniformly at random. The expected squared norm of stochastic gradients is uniformly bounded, i.e., $\left\|\nabla F_{i}\left(\mathbf{w}_{t}^{i},\xi_{j}\right)\right\|\leq G$ for all $i\in[1,\cdots,N]$ , $t\in[0,\cdots,T-1]$ and $j\in[1,\cdots,|\mathcal{B}_{i}|]$
Theorem 1
@@ VI-A Convergence Analysis
The meaning of $\Gamma$ is consistent with [29], where a larger value of $\Gamma$ indicates that the data among different clients is more Non-IID.
proof sketch: We investigate the relationship between $\mathbb{E}\|\mathbf{w}_{t+1}-\mathbf{w}^{*}\|^{2}$ and $\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$, and then use mathematical induction to obtain an upper bound for $\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$. Finally, by utilizing Assumption 1, we derive Equation (16). For the detailed proof, please refer to Appendix A-A.
VI-B Optimal Fairness Parameter $\lambda^{*}$
Assuming $Q_{0}\leq\Delta_{i}^{j}\leq Q_{1}$, then $1+Q_{0}\lambda\leq C_{t}^{i,j}\leq 1+Q_{1}\lambda$. We rearrange the convergence upper bound in Equation (16) to obtain a function in terms of $\lambda$,